Mar 3, 2021 Calorimeter
Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time
- To join via a Web Browser, go to the page  https://bluejeans.com/907185247.
- To join via Polycom room system go to the IP Address: 220.127.116.11 (bjn.vc) and enter the meeting ID: 907185247.
- To join via phone, use one of the following numbers and the Conference ID: 907185247.
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
- More information on connecting to bluejeans is available.
Participant Direct Lines
- JLab Phone in CC L207 is 757-269-7084 (usual room)
- JLab Phone: in CC F326 is 757-269-6460
- Phone in the Regina Video-conference Suite is 306-585-4204
- Athens Phone: in Christina's office is 011-30-210-727-6947
- FCAL HDFCAL log book
- BCAL HDBCAL log book
- Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
Goals for Calorimetry Group
- Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
- Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
- Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.
- Run Preparations
- BCAL LED
- BCAL (offline)
- FCAL (offline)
- Action Items
- Any other business
Attending: Elton, Karthik, Mark, Sean, Varun, Zisis, Susan, Sasha, Matt, Rebecca, Igal, Chandra, Jon, Tolga
- GlueX Review (March 25-26): nothing is expected from CALWG for this.
- COMPCAL paper: Sasha has received feedback and a new draft will be available on DoDB soon.
- Sean mentioned that new trees are available for the photon reconstruction to be looked at by Churamani.
- Run Preparations
- FCAL: Chris is replacing resistors on the send batch of 100 bases, which will leave about 200 to go after these. Nick is checking the base communications. Matt said that repaired based have been sent from IU to JLAB as well as a testing device with gives the voltages off all pins.
- BCAL: no report.
- CCAL: Sasha reported that the large amplified on the CCAL bases causes linearity issues; such amplification is not needed PrimEx. A smaller one x3 will be used. Also, the LMS system is being refurbished with 1mm plastic fibers. Discussion is ongoing with Fernando to separate the HV from amplifier lines for FCAL-II, and Fernando's group will make a prototype; this will not be in time for CCAL opportunistic testing during the SRC run. A discussion started on replacing the CW bases for FCAL-II with a conventional divider. This will need an investment from the lab.
- BCAL: Karthik reported that 15% of the z-calibration jobs are failing due to the existing tape/last files issues.
- BCAL LED: no report.
- BCAL (offline): Sean mentioned that the summer 2020 data will soon be available and Tolga will carry out his standard monitoring checks.
- FCAL (offline): Susan is also hampered by the tape business. Sean will direct here to the available runs on cache and Susan will continue with what is available. Susan will look at LED tigers and pulses from those.
- Reconstruction: Igal asked when the new batches will be ready for offsite production. Sean answered "after full validation", so in about 1 month from now. Igal asked to use IU Big Red for tests and Matt said that this is ok.
- Simulations: Sean asked about FCAL threshold handling, which is of interest to Daniel's work (MC resolutions). Matt responded that individual thresholds had been tuned using G4 in the past. Rebecca will circulate her slides on block-level thresholds (from raw hits). The focus was on the input to the neural bet at that time, which is not directly related to thresholds. Susan offered to help.
- Action Items:
- Short term: we reviewed those and assigned names to each task. Item 2 relates to both FCAL hole and FCAL-BCAL gap. Rebecca is looking into some studies. This information is useful for the physics analyses. Matt asked Mark if he tested the stability of the SDMEs under variation of these fiducial cuts. Mark will make the cuts tighter (using existing trees) and look at this because opening the cuts needs to be done at the analysis tree level.
- Long term: the list stands. Sean comments that the G3 default process behaviour is not seen in data.
- Work Packages: tweaks were made to these. Mark added that he could work with Susan. The reaction in the past was limited by kinematics in FCAL with a subset of the data using eta-pi, where the pi+pi- helps with the vertex. The assumption is that the eta position resolutions contributes little, and that the energy resolution dominates. They will discuss it offline and plan to use the full data set.
- Any other business